Gray and Juvia casually vibe with each other at the bottom of a 42 m deep lake. It is 2.1°C where they currently are, and 23 °C at the surface. Find the ratio of the volume of one of Juvia’s air bubbles at her current depth to its volume when it reaches the surface.”.


1
Expert's answer
2022-02-19T04:56:01-0500

Given:

"h=42\:\rm m"

"T_1=2.1+273=275.1\:\rm K"

"T_2=23+273=298\:\rm K"

"p_a=10^5\:\rm Pa"


The equation of state of a perfect gas says

"\frac{p_1V_1}{T_1}=\frac{p_2V_2}{T_2}"

Here

"p_1=p_a+g\rho h,\quad p_2=p_a"

The ratio of the volume of one of Juvia’s air bubbles at her current depth to its volume when it reaches the surface

"\frac{V_1}{V_2}=\frac{p_2T_1}{p_1T_2}=\frac{p_aT_1}{(p_a+g\rho h)T_2}"

"\frac{V_1}{V_2}=\frac{10^5*275.1}{(10^5+9.8*1000*45)*298}=0.17"